Cupcake123456:

Which expression is equivalent to 2(a+2b)-a-2b 3a+2b 3a-2b a-2b a+2b @razor @sourmunchkin7806

Razor:

2(a + 2b) - a - 2b = 2a + 4b - a - 2b = a + 2b

Razor:

soooooo a + 2b